A particle at opposition r=I+ 2j+ 3k is acted upon by a force F 2= 4i +5j +6k.


Evaluate the torque experienced by the particle.

1
Expert's answer
2022-03-29T12:27:58-0400

The torque experienced by the particle

"\vec \tau=\vec r\times \vec F"

Hence

"\vec \tau=(\hat i+ 2\hat j+ 3\hat k)\times (4\hat i +5\hat j +6\hat k)\\
=5\hat k-6\hat j-8\hat k+12 \hat i+12 \hat j-15 \hat i\\
=-3\hat i+6\hat j-3\hat k"